OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4. A.)2 B.) 3 C.) All real numbers D.) No solution

OpenStudy (texaschic101):

-2(x + 3) = -2(x + 1) - 4 --- distribute through the parenthesis -2x - 6 = -2x - 2 - 4 --- combine like terms -2x - 6 = -2x - 6 they are the same, therefore, there are infinite solutions. C -- all real numbers.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Can you help me with others

OpenStudy (texaschic101):

sure....what ya got ?

OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar. Answer for Blank 1:

OpenStudy (texaschic101):

(3/4)x + 5/8 = 4x --- is this your problem ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4. A.) 1 B.) -1 C.) -5 D.) -25

OpenStudy (texaschic101):

(3/4)x + 5/8 = 4x 5/8 = 4x - 3/4x 5/8 = 16/4x - 3/4x 5/8 = 13/4x (5/8) / (13/4) = x 5/8 * 4/13 = x 20/104 = x 5/26 = x check.. 3/4(5/26) + 5/8 = 4(5/26) 15/104 + 5/8 = 20/26 15/104 + 65/104 = 20/26 80/104 = 20/26 10/13 = 10/13 (correct) solution is : x = 5/26 =============== 3(x + 1) = -2(x - 1) - 4 3x + 3 = -2x + 2 - 4 3x + 3 = -2x - 2 3x + 2x = -2 - 3 5x = -5 x = -1

OpenStudy (texaschic101):

sorry it took so long..I was having a little trouble with the first one. Do you have any questions ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The triangle below has a total perimeter of 450 cm. picture of a triangle with sides 128 and 150. Third side unknown Write an equation and solve it to find the length of the longest side of the triangle. A.)128x + 150 = 450; x = 2.34 cm B.) x + 278 = 450; x = 172 cm C.) 2x + 278 = 450; x = 86 cm D.) 150x + 128 = 450; x = 2.15 cm

OpenStudy (anonymous):

Solve for x: the quantity of x plus 16 all over 3 = 3x A.) x = -7 B.) x = -13 C.) x = 8 D.) x = 2

OpenStudy (texaschic101):

P = A + B + C 450 = 128 + 150 + C 450 = 278 + C 450 - 278 = C 172 = C ANSWER : B ============ (x + 16)/3 = 3x --- multiply by 3 x + 16 = 9x 16 = 9x - x 16 = 8x 16/8 = x 2 = x ANSWER : D

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

no problem :)